### v2

@@ Abstract
In nonperturbative formulation of Euclidean signature quantum field theory (QFT), the vacuum state is characterized by the Wilsonian renormalization group (RG) flow of Feynman measures. Such an RG flow is a family of Feynman measures on the space of ultraviolet (UV) regularized fields, linked by the Wilsonian renormalization group equation. In this paper we show that under mild conditions, a Wilsonian RG flow of Feynman measures extending to arbitrary regularization strenghts has a factorization property: there exists an ultimate Feynman measure (UV limit) on the distribution sense fields, such that the regularized instances in the flow are obtained from this UV limit via taking the marginal measure against the regulator. Existence theorems on the flow and UV limit of the corresponding action functional are also discussed.
keywords:
Wilsonian renormalization, renormalization group flow, UV limit
@@ 1 Introduction
It is well known, that a Gaussian Feynman measure $\gamma$ on the space of distribution sense fields can be naturally associated to the kinetic term $T$, as follows [1, 2, 3, 4, 5]. Denoting by $\mathtt{K}$ the maximally symmetric fundamental solution (Green’s distribution) to the partial differential operator $(-\Delta+m^{2})$ subordinate to $T$, the Gaussian function $Z_{\mathrm{free}}(j):=\mathrm{e}^{-\frac{1}{2}(j\,|\,\mathtt{K}\star j)}$ is the Fourier transform of a probability measure $\gamma$ on the space of distribution sense fields, as implied by the well known Bochner–Minlos theorem (see e.g. [2] Corollary 1 for a concise review). This Gaussian measure $\gamma$ is a genuine probability measure, devoid of any pathologies, and is the mathematically precise definition of the Feynman measure for a non-interacting Euclidean signature QFT model. It is customary to write $\int(\dots)\,\mathrm{d}\gamma(\phi)$ informally as $\int(\dots)\,\mathrm{e}^{-T(\phi)}\,\mathrm{d}\phi$, the integrals performed on the distribution sense fields.
The tentative definition for the Feynman measure of the interacting QFT model is the product measure $\mu:=\mathrm{e}^{-V}\cdot\gamma$, meaning that $\int(\dots)\,\mathrm{d}\mu(\phi):=\int(\dots)\,\mathrm{e}^{-V(\phi)}\,\mathrm{d}\gamma(\phi)$, where the right hand side of the equation is the tentative rigorous definition for the informal expression $\int(\dots)\,\mathrm{e}^{-(T+V)(\phi)}\,\mathrm{d}\phi$. This construction is known to be rather obviously problematic: the measure $\gamma$ naturally lives on the space of distribution sense fields, whereas the Feynman exponent $\mathrm{e}^{-V}$ of interaction term lives on the space of function sense fields. Thus, the product measure $e^{-V}\cdot\gamma$ cannot be naively defined.
@@
To see that the above concept indeed corresponds to the usual constructive implementation of the Wilsonian RG, consider a free massive Klein–Gordon QFT encoded by its Gaussian measure $\gamma$ on the space of tempered distributions. For every $\Lambda>0$, consider the linear operator $C_{\Lambda}$ which cuts off the momentum modes $p^{2}>\Lambda^{2}$ of a tempered distribution, whenever that operation is meaningful. (This amounts to the multiplication of a tempered distribution with a non-smooth function which is not everywhere defined.) One can show that $C_{\Lambda}$ is indeed a well defined $\gamma$-measurable mapping in a certain sense, and that the pushfoward measures $(C_{\Lambda})_{*}\,\gamma$ indeed correspond to the usual UV-cutoff versions of $\gamma$. The next step is to include a family of Feynman exponents $e^{-V_{\Lambda}}$ ($0<\Lambda<\infty$) of interaction terms in a way that implements Wilson’s idea: namely to integrate out high-momentum modes in every step. This is precisely what the momentum cutoff operators do: the RG operation connecting $\mathrm{e}^{-V_{\Lambda^{\prime}}}\cdot(C_{\Lambda^{\prime}})_{*}\,\gamma$ with $\mathrm{e}^{-V_{\Lambda}}\cdot(C_{\Lambda})_{*}\,\gamma$ ($0<\Lambda^{\prime}<\Lambda<\infty$) corresponds to the pushforward by $C_{\Lambda^{\prime}}$, modulo an eventual field rescaling by factors $z(\Lambda^{\prime})$ and $z(\Lambda)$.
Because the measurability of the sharp UV cutoff operators is intimately tied to the free theory, we will focus on the more generally applicable approach of using continuous coarse-graining operators (practically corresponding to smooth momentum cutoffs). The benefit is that the corresponding pushforward operations are well-defined irrespective of which Borel measure they act on such that the setup generalises to arbitrary scalar quantum field theories on either flat or curved spacetime, and without an a priori splitting of the model to a free plus an interaction term. The precise form of the RG equation reads
@@
In this paper we prove structural theorems on the space of nonterminating Wilsonian RG flows. We show that for scalar fields over a flat spacetime manifold, given a Wilsonian RG flow of measures $\mu_{C}$ for $C$ in a suitable class of coarse-grainings, there exists a unique measure $\mu$ from which the flow originates. Moreover, we prove that if two Wilsonian RG flows $\mu_{C}$ and $\gamma_{C}$ have an interaction potential $V_{C}$ relative to each-other such that $\mu_{C}=\mathrm{e}^{-V_{C}}\cdot\gamma_{C}$ holds for a suitably chosen coarse-gaining $C$, then a potential $V_{C}$ will exist for all coarse-gainings $C$, and the flow of potentials will have a UV limit $V$ on the distributional sense fields such that $\mu=\mathrm{e}^{-V}\cdot\gamma$ holds between the UV limit measures. It will also be shown that $V$ has the same lower bound as $V_{C}$ if those potentials were bounded from below at a suitably chosen coarse-graining $C$. We also point out possibilities for strengthening the mentioned results to manifold spacetimes.
The above result can be seen as the analogy of the existence theorem for the UV limit of Wilsonian RG flows of Feynman correlators in arbitrary signatures, proved in [16]. In the present paper, however, due to the Euclidean signature setting, more constraining statements can be proved. For rigorous studies on RG flows of Euclidean signature theories we also refer to the construction in [17], which formalizes the notion of Wetterich type RG flows. We note, however, that the notion of Wilsonian and the Wetterich RG flow constructions are not literally the same, moreover that rigorous theorems on the UV behavior of RG flows are generally rather scarce in the literature.
@@
The structure of the paper is as follows. In Section 2 the mathematical framework for Wilsonian RG flows of regularized Feynman measures is rigorously spelled out. In Section 3 we prove the existence theorem for a UV limit Feynman measure. In Section 4 we prove an existence theorem for the UV limit of the flow of relative interaction potentials between two flows. Based on the presented theorems, in Section 5 case studies are presented on concrete QFT models applying the above results. In Section 6 a conclusion is presented. The paper is closed by Appendix A, a review on some known important properties of Gaussian measures on the space of distribution sense fields.
@@
The following normalization convention should be noted. The running field renormalization factor appearing in (1) may be merged into the measures $\mu_{C}$ in the flow, by considering rather the flow of measures $z(C)_{*}\,\mu_{C}$ on the space of rescaled fields ($C\in\{\mbox{coarse-grainings}\}$}). That convention, in QFT terms, corresponds to the book-keeping of the running field renormalization factor as the running coupling of the kinetic term in the action functional. This convention sets $z(C)=1$ in (1), which is used from this point on, and therefore the running field renormalization factor will not appear in the notation explicitely in the sequel, although it is accounted for. In the pertinent convention, the factor re-appears when comparing two flows.
@@ 2 Wilsonian renormalization group (RG) flows
Throughout the paper, the standard distribution theory notations are used [18, 19]: $\mathcal{E}$ for $\mathbb{R}^{N}\to\mathbb{C}$ smooth functions, $\mathcal{S}$ for Schwartz functions, $\mathcal{D}$ for test functions, understood with their standard topologies, and $\mathcal{E}^{\prime}$, $\mathcal{S}^{\prime}$, $\mathcal{D}^{\prime}$ for their continuous duals, understood with their standard strong dual topologies. In particular, $\mathcal{E}^{\prime}$ is the space of compactly supported distributions, $\mathcal{S}^{\prime}$ is the space of tempered distributions, and $\mathcal{D}^{\prime}$ is the space of all distributions. Note that the spaces $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are only meaningful on flat spacetimes, whereas the spaces $\mathcal{E}$, $\mathcal{D}$ and $\mathcal{E}^{\prime}$, $\mathcal{D}^{\prime}$ are also meaningful on any orientable and oriented manifold without further assumptions. The symbol $F$ shall denote the usual $F:\,\mathcal{S}\to\mathcal{S}$ Fourier transform, along with its extension $F:\,\mathcal{S}^{\prime}\to\mathcal{S}^{\prime}$ to the space of tempered distributions. Both of these are known to be topological vector space automorphisms. The normalization convention of $F$ is such that for all $\varphi,\psi\in\mathcal{S}$ the identity $F(\varphi\star\psi)=F(\varphi)\cdot F(\psi)$ holds, where $\star$ denotes convolution and $\cdot$ denotes pointwise product. The above topological vector spaces $\mathcal{E}$, $\mathcal{S}$, $\mathcal{D}$, and $\mathcal{E}^{\prime}$, $\mathcal{S}^{\prime}$, $\mathcal{D}^{\prime}$ are understood together with the Borel sets generated by their standard topology. For an $\eta\in\mathcal{S}$ we equip the subspace $\eta\star\mathcal{S}^{\prime}\subset\mathcal{S}^{\prime}$ with the subspace topology, and corresponding Borel sets. Often, we denote the convolution operator by an $\eta\in\mathcal{S}$ using the notation $C_{\eta}:\,\mathcal{S}\to\mathcal{S}$ and $C_{\eta}:\,\mathcal{S}^{\prime}\to\mathcal{S}^{\prime}$, furthermore we use the notation $C_{\eta}\,t=\eta\star t$ given some $t\in\mathcal{S}^{\prime}$ interchangeably, according to convenience. In particular, one has $C_{\eta}[\mathcal{S}^{\prime}]=\eta\star\mathcal{S}^{\prime}$ by notation. The standard symbols $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ and $\overline{\mathbb{R}}_{0}^{+}:=\mathbb{R}_{0}^{+}\cup\{+\infty\}$ are used for the extended real numbers and extended non-negative real numbers, along with $\mathrm{e}^{-\infty}:=0$ and $\mathrm{e}^{+\infty}:=+\infty$, furthermore we use the convention $\pm\infty\cdot 0:=0$ as customary in measure theory.
@@
holds. Here, $()_{*}$ denotes the pushforward of a measure by a measurable map, i.e. by definition $\big{(}(C_{\eta^{\prime}})_{*}\,\mu_{\eta}\big{)}(B):=\mu_{\eta}\big{(}\overset{-1}{C_{\eta^{\prime}}}(B)\big{)}$ for all Borel sets $B$, where $\overset{-1}{C_{\eta^{\prime}}}(B)$ denotes the preimage of $B$ by the mapping $C_{\eta^{\prime}}$.
@@
    It is well known, but [19] Proposition A.9 and after explicetely shows that the function spaces $\mathcal{E}$, $\mathcal{S}$, $\mathcal{D}$ and their strong duals $\mathcal{E}^{\prime}$, $\mathcal{S}^{\prime}$, $\mathcal{D}^{\prime}$ are Souslin spaces.
@@ 3 Existence theorem of the UV limit measure
    In addition, [29] and the proof of [30] Lemma 1 shows that in the above factorization, one of the factors, say $\eta$, may be chosen to be $F(\eta)>0$. Quite obviously, the normalization of such $\eta$ may be chosen ensuring that $F(\eta)$ is unity at the origin. Due to the paracompactness of $\mathbb{R}^{N}$, for such $\eta$ one may construct a compactly supported smooth function $\alpha:\,\mathbb{R}^{N}\to[0,1]$ which is unity over a compact neighborhood of the origin containing the level set $\{F(\eta)\geq\frac{1}{2}\}$. Clearly, the function $\Phi:=(1-\alpha)\cdot 1+\alpha\cdot 1/F(\eta)$ is smooth, everywhere positive, and is unity outside the support of $\alpha$, moreover $F^{-1}(\Phi\cdot F(\eta))\in\mathcal{R}_{\mathcal{S}+}$, see again Definition 1. Furthermore, for any $\ell\in\mathcal{S}$ one has $F^{-1}(1/\Phi\cdot F(\ell))\in\mathcal{S}$. Therefore, without loss of generality, the $\eta$ in (i) may be chosen to be $\eta\in\mathcal{R}_{\mathcal{S}+}$.
@@
(factorizarion) For every $j\in\mathcal{S}$ there exist $\eta\in\mathcal{S}$ and $\ell\in\mathcal{S}$, such that $j=\eta\star\ell$ holds. (This $\eta$ exists also such that $\eta\in\mathcal{R}_{\mathcal{S}}$ or even $\eta\in\mathcal{R}_{\mathcal{S}+}$.)
@@
Let $\eta\in\mathcal{S}$ with nowhere vanishing $F(\eta)$. Then, the convolution operator $C_{\eta}:\,\mathcal{S}^{\prime}\to\mathcal{S}^{\prime},\,t\mapsto C_{\eta}\,t:=\eta\star t$ is injective. In particular, then $C_{\eta}:\,\mathcal{S}\to\mathcal{S}$ is injective.
@@
Take some $\eta\in\mathcal{S}$ and $t\in\mathcal{S}^{\prime}$. Then the Fourier transform of the convolution of $\eta$ and $t$ is $F(\eta\star t)=F(t)F(\eta)$ as tempered distribution, meaning that for all $\varphi\in\mathcal{S}$ one has $(F(\eta\star t)\,|\,\varphi)=(F(t)\,|\,F(\eta)\cdot\varphi)$. In particular, for all $\varphi\in\mathcal{D}\subset\mathcal{S}$ that identity holds. Consider now such an $\eta$ which has nowhere vanishing $F(\eta)$. Then for any $\varphi\in\mathcal{D}$ the function $\varphi/F(\eta)$ is everywhere defined, smooth and compactly supported, i.e. $\varphi/F(\eta)\in\mathcal{D}$. Consequently, for all $\varphi\in\mathcal{D}$ one has $(F(\eta\star t)\,|\,\varphi/F(\eta))=(F(t)\,|\,\varphi)$. If $t\in\mathcal{S}^{\prime}$ is such that $\eta\star t=0$ holds, then by means of our observation, for all $\varphi\in\mathcal{D}$ the identity $(F(t)\,|\,\varphi)=(F(\eta\star t)\,|\,\varphi/F(\eta))=0$ needs to hold. That is, the tempered distribution $F(t)\in\mathcal{S}^{\prime}$ is zero on the dense subspace $\mathcal{D}\subset\mathcal{S}$, implying that $F(t)=0$. By the injectivity of the Fourier transformation, $t=0$ follows. ∎
@@
    It is common knowledge that the space $\mathcal{S}$ also admits approximate identity in the following sense. Sequences $\alpha_{n}\in\mathcal{R}_{\mathcal{S}}\subset\mathcal{S}$ ($n\in\mathbb{N}$) exist such that for all $\varphi\in\mathcal{S}$ one has $\alpha_{n}\star\varphi\to\varphi$ in the $\mathcal{S}$ topology as $n\to\infty$. For let us take any $\alpha\in\mathcal{R}_{\mathcal{S}}$ and define its compressed version $\alpha_{\Lambda}(x):=\Lambda^{N}\,\alpha(\Lambda\,x)$ ($x\in\mathbb{R}^{N}$) for all $1\leq\Lambda<\infty$. One can verify that for all $\varphi\in\mathcal{S}$ one has $\alpha_{\Lambda}\star\varphi\to\varphi$ in the $\mathcal{S}$ topology as $\Lambda\to\infty$. To confirm this, one should note that $F:\,\mathcal{S}\to\mathcal{S}$ is topological isomorphism, and therefore it is enough to show that $(1-F(\alpha_{\Lambda}))\cdot F(\varphi)\to 0$ in the $\mathcal{S}$ topology as $\Lambda\to\infty$. The canonical topology of $\mathcal{S}$ is characterized by the family of seminorms $\sup_{x\in\mathbb{R}^{N}}|Q(x)\,P(\partial_{x})\,\psi(x)|$ on a $\psi\in\mathcal{S}$, with $P,Q$ being any polynomials. Clearly, the seminorms $\sup_{x\in\mathbb{R}^{N}}|P(\partial_{x})\,Q(x)\,\psi(x)|$ ($\psi\in\mathcal{S}$) define the very same topology. Due to Hölder’s inequality and the Sobolev inequality, this family of seminorms is equivalent to the familiy of seminorms $\|P(\partial)\,Q\,\psi\|_{L^{2}}$ ($\psi\in\mathcal{S}$), with $P,Q$ being polynomials. Using $\left.F(\alpha_{\Lambda})\right|_{\omega}=\left.F(\alpha)\right|_{\omega/\Lambda}$ ($\omega\in\mathbb{R}^{N}$), via Lebesgue’s theorem of dominated convergence it is not hard to see that for all polynomials $P,Q$ and all $\alpha\in\mathcal{R}_{\mathcal{S}}$, $\varphi\in\mathcal{S}$ one has $\|P(\partial)\,Q\cdot(1-F(\alpha_{\Lambda}))\cdot F(\varphi)\|_{L^{2}}\to 0$ as $\Lambda\to\infty$, which completes the argument.
@@
$Z(j)$ $:=$ $\int\limits_{\phi\in\mathcal{S}^{\prime}}\mathrm{e}^{\mathrm{i}\,(\phi|j)}\,\mathrm{d}\mu(\phi)\qquad\mbox{ for all }j\in\mathcal{S}.$ (11)
@@
$Z(j_{n})=Z(\eta\star\ell_{n})=Z_{\eta^{t}}(\ell_{n})\;\to\;Z_{\eta^{t}}(\ell)=Z(\eta\star\ell)=Z(j)\qquad(\mbox{as }n\to\infty),$ (18)

### v3

@@ Abstract
In nonperturbative formulation of Euclidean signature quantum field theory (QFT), the vacuum state is characterized by the Wilsonian renormalization group (RG) flow of Feynman measures. Such an RG flow is a family of Feynman measures on the space of ultraviolet (UV) regularized fields, linked by the Wilsonian renormalization group equation. In this paper we show that under mild conditions, a Wilsonian RG flow of Feynman measures extending to arbitrary regularization strengths has a factorization property: there exists an ultimate Feynman measure (UV limit) on the distribution sense fields, such that the regularized instances in the flow are obtained from this UV limit via taking the marginal measure against the regulator. Existence theorems on the flow and UV limit of the corresponding action functional are also discussed.
Keywords: Wilsonian renormalization, renormalization group flow, UV limit
@@ 1 Introduction
It is well known, that a Gaussian Feynman measure $\gamma$ on the space of distribution sense fields can be naturally associated to the kinetic term $T$, as follows [1, 2, 3, 4, 5]. Denoting by $\mathtt{K}$ the maximally symmetric fundamental solution (Green’s distribution) to the partial differential operator $(-\Delta+m^{2})$ subordinate to $T$, the Gaussian function $Z_{\mathrm{free}}(j):=\rme^{-\frac{1}{2}(j\,|\,\mathtt{K}\star j)}$ is the Fourier transform of a probability measure $\gamma$ on the space of distribution sense fields, as implied by the well known Bochner–Minlos theorem (see e.g. [2] Corollary 1 for a concise review). This Gaussian measure $\gamma$ is a genuine probability measure, devoid of any pathologies, and is the mathematically precise definition of the Feynman measure for a non-interacting Euclidean signature QFT model. It is customary to write $\int(\dots)\,\mathrm{d}\gamma(\phi)$ informally as $\int(\dots)\,\rme^{-T(\phi)}\,\mathrm{d}\phi$, the integrals performed on the distribution sense fields.
The tentative definition for the Feynman measure of the interacting QFT model is the product measure $\mu:=\rme^{-V}\cdot\gamma$, meaning that $\int(\dots)\,\mathrm{d}\mu(\phi):=\int(\dots)\,\rme^{-V(\phi)}\,\mathrm{d}\gamma(\phi)$, where the right hand side of the equation is the tentative rigorous definition for the informal expression $\int(\dots)\,\rme^{-(T+V)(\phi)}\,\mathrm{d}\phi$. This construction is known to be rather obviously problematic: the measure $\gamma$ naturally lives on the space of distribution sense fields, whereas the Feynman exponent $\rme^{-V}$ of interaction term lives on the space of function sense fields. Thus, the product measure $e^{-V}\cdot\gamma$ cannot be naively defined.
@@
To see that the above concept indeed corresponds to the usual constructive implementation of the Wilsonian RG, consider a free massive Klein–Gordon QFT encoded by its Gaussian measure $\gamma$ on the space of tempered distributions. For every $\Lambda>0$, consider the linear operator $C_{\Lambda}$ which cuts off the momentum modes $p^{2}>\Lambda^{2}$ of a tempered distribution, whenever that operation is meaningful. (This amounts to the multiplication of a tempered distribution with a non-smooth function which is not everywhere defined.) One can show that $C_{\Lambda}$ is indeed a well defined $\gamma$-measurable mapping in a certain sense, and that the pushforward measures $(C_{\Lambda})_{*}\,\gamma$ indeed correspond to the usual UV-cutoff versions of $\gamma$. The next step is to include a family of Feynman exponents $e^{-V_{\Lambda}}$ ($0<\Lambda<\infty$) of interaction terms in a way that implements Wilson’s idea: namely to integrate out high-momentum modes in every step. This is precisely what the momentum cutoff operators do: the RG operation connecting $\rme^{-V_{\Lambda^{\prime}}}\cdot(C_{\Lambda^{\prime}})_{*}\,\gamma$ with $\rme^{-V_{\Lambda}}\cdot(C_{\Lambda})_{*}\,\gamma$ ($0<\Lambda^{\prime}<\Lambda<\infty$) corresponds to the pushforward by $C_{\Lambda^{\prime}}$, modulo an eventual field rescaling by factors $z(\Lambda^{\prime})$ and $z(\Lambda)$.
Because the measurability of the sharp UV cutoff operators is intimately tied to the free theory, we will focus on the more generally applicable approach of using continuous coarse-graining operators (practically corresponding to smooth momentum cutoffs). The benefit is that the corresponding pushforward operations are well-defined irrespective of which Borel measure they act on such that the setup generalizes to arbitrary scalar quantum field theories on either flat or curved spacetime, and without an a priori splitting of the model to a free plus an interaction term. The precise form of the RG equation reads
@@
In this paper we prove structural theorems on the space of nonterminating Wilsonian RG flows. We show that for scalar fields over a flat spacetime manifold, given a Wilsonian RG flow of measures $\mu_{C}$ for $C$ in a suitable class of coarse-grainings, there exists a unique measure $\mu$ from which the flow originates. Moreover, we prove that if two Wilsonian RG flows $\mu_{C}$ and $\gamma_{C}$ have an interaction potential $V_{C}$ relative to each-other such that $\mu_{C}=\rme^{-V_{C}}\cdot\gamma_{C}$ holds for a suitably chosen coarse-gaining $C$, then a potential $V_{C}$ will exist for all coarse-gainings $C$, and the flow of potentials will have a UV limit $V$ on the distributional sense fields such that $\mu=\rme^{-V}\cdot\gamma$ holds between the UV limit measures. It will also be shown that $V$ has the same lower bound as $V_{C}$ if those potentials were bounded from below at a suitably chosen coarse-graining $C$. We also point out possibilities for strengthening the mentioned results to manifold spacetimes. The UV measure existence result can be seen as the analogy of the existence theorem for the UV limit of Wilsonian RG flows of Feynman correlators in arbitrary signatures, proved in [16]. In the present paper, however, due to the Euclidean signature setting, more constraining statements can be proved.
The shown UV theorems on nonterminating Wilsonian RG flows are also discussed in the light of certain concrete models, such as the $\varphi^{4}$ model in various dimensions, and on models with bounded potential density. With the help of the presented theorems one can demonstrate, for instance, that the limiting procedure usually applied for constructing Feynman measures of high dimensional $\varphi^{4}$ theories [17, 18, 19, 20] are themselves non-Wilsonian families. A renormalizability result on bounded potential density is also presented.
There is a difference, when comparing the described Wilsonian RG prescription to other constructive approaches in QFT, as considered e.g. in [21]. There, one consideres an already regularized instance of the model in question, living on a lattice in a finite volume. Then, one uses sophisticated techniques such as polymer expansions to obtain bounds on partition functions or Schwinger functions which will inevitably depend on the volume as well as the lattice spacing. The next step is to show that these bounds are sufficiently uniform to take the continuum limit. Often, this is established utilizing the Wilsonian RG as a technique to split the integration over all fields into integration over fields with their momenta in specific intervals (momentum slices, see e.g. [22]). The end result is then a family of measures $(\mu_{\Lambda})$ indexed by UV-cutoffs $\Lambda$ which converges to some UV limit $\mu$. In other similar, but non-lattice constructions [18, 19] one works on the genuine continuum Euclidean spacetime, but with UV regularized fields to define a family $(\mu_{\Lambda})$ convergent to some UV measure $\mu$. It needs to be stressed that in all these prescriptions, the family $(\mu_{\Lambda})$ will in general not be Wilsonian itself, in the sense that it does not fulfill Wilsonian RG equation (1), as will be shown by our results. What is true, on the other hand, that in certain cases, they do converge to a UV measure $\mu$ in the so-called weak topology. Given this UV limit $\mu$, it is therefore a natural open question to ask, what the Wilsonian RG flow induced by the UV limit $\mu$ looks like, even though $\mu$ itself may have been defined as the limit of a non-Wilsonian regularized family. The non-Wilsonianity will be discussed in Example 35 in the context of $\varphi^{4}$ models.
For rigorous studies on RG flows of Euclidean signature theories we also refer to the construction in [23], which formalizes the notion of Wetterich type RG flows. While the underlying idea, to suppress high-momentum modes inside a functional integral, remains the same, it is not implemented via pushforwards (neither sharp nor continuous). Instead, one modifies the interaction potential $V_{\Lambda}$ by adding a bilinear term $B_{k}$ that suppresses momenta below the auxiliary scale $k\geq 0$. Furthermore, the resulting flow deals with the effective average actions, which are Legendre transforms of cumulant-generating functions, instead of the flow of the measures themselves.
@@
The structure of the paper is as follows. In Section 2 the mathematical framework for Wilsonian RG flows of regularized Feynman measures is rigorously spelled out. In Section 3 we prove the existence theorem for a UV limit Feynman measure. In Section 4 we prove an existence theorem for the UV limit of the flow of relative interaction potentials between two flows. Based on the presented theorems, in Section 5 case studies are presented on concrete QFT models applying the above results. In Section 6 a conclusion is presented. The paper is closed by A, a review on some known important properties of Gaussian measures on the space of distribution sense fields.
@@
The following normalization convention should be noted. The running field renormalization factor appearing in (1) may be merged into the measures $\mu_{C}$ in the flow, by considering rather the flow of measures $z(C)_{*}\,\mu_{C}$ on the space of rescaled fields ($C\in\{\mbox{coarse-grainings}\}$}). That convention, in QFT terms, corresponds to the book-keeping of the running field renormalization factor as the running coupling of the kinetic term in the action functional. This convention sets $z(C)=1$ in (1), which is used from this point on, and therefore the running field renormalization factor will not appear in the notation explicitly in the sequel, although it is accounted for. In the pertinent convention, the factor re-appears when comparing two flows.
@@ 2 Wilsonian renormalization group (RG) flows
[…]
@@
[…]
@@
    It is well known, but [25] Proposition A.9 and after explicitly shows that the function spaces $\mathcal{E}$, $\mathcal{S}$, $\mathcal{D}$ and their strong duals $\mathcal{E}^{\prime}$, $\mathcal{S}^{\prime}$, $\mathcal{D}^{\prime}$ are Souslin spaces.
@@ 3 Existence theorem of the UV limit measure
    In addition, [33] and the proof of [34] Lemma 1 shows that in the above factorization, one of the factors, say $\eta$, may be chosen to be $F(\eta)>0$. Quite obviously, the normalization of such $\eta$ may be chosen ensuring that $F(\eta)$ is unity at the origin. Due to the paracompactness of $\mathbb{R}^{N}$, for such $\eta$ one may construct a compactly supported smooth symmetric function $\alpha:\,\mathbb{R}^{N}\to[0,1]$ which is unity over a compact neighborhood of the origin containing the level set $\{F(\eta)\geq\frac{1}{2}\}$. Clearly, the function $\Phi:=(1-\alpha)\cdot 1+\alpha\cdot 1/F(\eta)$ is smooth, symmetric, everywhere positive, and is unity outside the support of $\alpha$, moreover $F^{-1}(\Phi\cdot F(\eta))\in\mathcal{R}_{\mathcal{S}+}$, see again Definition 1. Furthermore, for any $\ell\in\mathcal{S}$ one has $F^{-1}(1/\Phi\cdot F(\ell))\in\mathcal{S}$. Therefore, without loss of generality, the $\eta$ in (i) may be chosen to be $\eta\in\mathcal{R}_{\mathcal{S}+}$.
@@
(factorization) For every $j\in\mathcal{S}$ there exist $\eta\in\mathcal{S}$ and $\ell\in\mathcal{S}$, such that $j=\eta\star\ell$ holds. (This $\eta$ exists also such that $\eta\in\mathcal{R}_{\mathcal{S}}$ or even $\eta\in\mathcal{R}_{\mathcal{S}+}$.)
@@
Let $\eta\in\mathcal{S}_{\mathbb{C}}$ with nowhere vanishing $F(\eta)$. Then, the convolution operator $C_{\eta}:\,\mathcal{S}_{\mathbb{C}}^{\prime}\to\mathcal{S}_{\mathbb{C}}^{\prime},\,t\mapsto C_{\eta}\,t:=\eta\star t$ is injective. In particular, then $C_{\eta}:\,\mathcal{S}_{\mathbb{C}}\to\mathcal{S}_{\mathbb{C}}$ is injective.
@@
Take some $\eta\in\mathcal{S}_{\mathbb{C}}$ and $t\in\mathcal{S}_{\mathbb{C}}^{\prime}$. Then the Fourier transform of the convolution of $\eta$ and $t$ is $F(\eta\star t)=F(t)F(\eta)$ as tempered distribution, meaning that for all $\varphi\in\mathcal{S}_{\mathbb{C}}$ one has $(F(\eta\star t)\,|\,\varphi)=(F(t)\,|\,F(\eta)\cdot\varphi)$. In particular, for all $\varphi\in\mathcal{D}_{\mathbb{C}}\subset\mathcal{S}_{\mathbb{C}}$ that identity holds. Consider now such an $\eta$ which has nowhere vanishing $F(\eta)$. Then for any $\varphi\in\mathcal{D}_{\mathbb{C}}$ the function $\varphi/F(\eta)$ is everywhere defined, smooth and compactly supported, i.e. $\varphi/F(\eta)\in\mathcal{D}_{\mathbb{C}}$. Consequently, for all $\varphi\in\mathcal{D}_{\mathbb{C}}$ one has $(F(\eta\star t)\,|\,\varphi/F(\eta))=(F(t)\,|\,\varphi)$. If $t\in\mathcal{S}^{\prime}$ is such that $\eta\star t=0$ holds, then by means of our observation, for all $\varphi\in\mathcal{D}_{\mathbb{C}}$ the identity $(F(t)\,|\,\varphi)=(F(\eta\star t)\,|\,\varphi/F(\eta))=0$ needs to hold. That is, the tempered distribution $F(t)\in\mathcal{S}_{\mathbb{C}}^{\prime}$ is zero on the dense subspace $\mathcal{D}_{\mathbb{C}}\subset\mathcal{S}_{\mathbb{C}}$, implying that $F(t)=0$. By the injectivity of the Fourier transformation, $t=0$ follows. ∎
@@
    It is common knowledge that the space $\mathcal{S}$ also admits approximate identity in the following sense. Sequences $\alpha_{n}\in\mathcal{R}_{\mathcal{S}}\subset\mathcal{S}$ ($n\in\mathbb{N}$) exist such that for all $\varphi\in\mathcal{S}$ one has $\alpha_{n}\star\varphi\to\varphi$ in the $\mathcal{S}$ topology as $n\to\infty$. For let us take any $\alpha\in\mathcal{R}_{\mathcal{S}}$ and define its compressed version $\alpha_{\Lambda}(x):=\Lambda^{N}\,\alpha(\Lambda\,x)$ ($x\in\mathbb{R}^{N}$) for all $1\leq\Lambda<\infty$. One can verify that for all $\varphi\in\mathcal{S}$ one has $\alpha_{\Lambda}\star\varphi\to\varphi$ in the $\mathcal{S}$ topology as $\Lambda\to\infty$. To confirm this, one should note that $F:\,\mathcal{S}_{\mathbb{C}}\to\mathcal{S}_{\mathbb{C}}$ is topological automorphism, and therefore it is enough to show that $(1-F(\alpha_{\Lambda}))\cdot F(\varphi)\to 0$ in the $\mathcal{S}$ topology as $\Lambda\to\infty$. The canonical topology of $\mathcal{S}$ is characterized by the family of seminorms $\sup_{x\in\mathbb{R}^{N}}|Q(x)\,P(\partial_{x})\,\psi(x)|$ on a $\psi\in\mathcal{S}_{\mathbb{C}}$, with $P,Q$ being any polynomials. Clearly, the seminorms $\sup_{x\in\mathbb{R}^{N}}|P(\partial_{x})\,Q(x)\,\psi(x)|$ ($\psi\in\mathcal{S}_{\mathbb{C}}$) define the very same topology. Due to Hölder’s inequality and the Sobolev inequality, this family of seminorms is equivalent to the family of seminorms $\|P(\partial)\,Q\,\psi\|_{L^{2}}$ ($\psi\in\mathcal{S}_{\mathbb{C}}$), with $P,Q$ being polynomials. Using $\left.F(\alpha_{\Lambda})\right|_{\omega}=\left.F(\alpha)\right|_{\omega/\Lambda}$ ($\omega\in\mathbb{R}^{N}$), via Lebesgue’s theorem of dominated convergence it is not hard to see that for all polynomials $P,Q$ and all $\alpha\in\mathcal{R}_{\mathcal{S}}$, $\varphi\in\mathcal{S}$ one has $\|P(\partial)\,Q\cdot(1-F(\alpha_{\Lambda}))\cdot F(\varphi)\|_{L^{2}}\to 0$ as $\Lambda\to\infty$, which completes the argument.
@@
$Z(j)$ $:=$ $\int\limits_{\phi\in\mathcal{S}^{\prime}}\rme^{\mathrm{i}\,(\phi|j)}\,\mathrm{d}\mu(\phi)\qquad\mbox{ for all }j\in\mathcal{S}.$ (12)
@@
$Z(j_{n})=Z(\eta\star\ell_{n})=Z_{\eta^{t}}(\ell_{n})\;\to\;Z_{\eta^{t}}(\ell)=Z(\eta\star\ell)=Z(j)\quad(\mbox{as }n\to\infty),$ (19)